OpenStudy (anonymous):

help me find the zero f(x)=(x-2)^2-9

OpenStudy (imstuck):

In order to find the zeros you have to factor it.

OpenStudy (anonymous):

\(\Large f(x)=(x-2)^2-9=(x-2+3)(x-2-3)=(x+1)(x-5)\)

OpenStudy (imstuck):

Expanding that you get\[f(x)=x ^{2}-4x+4-9\]

OpenStudy (imstuck):

Set it equal to 0 and get \[0=x ^{2}-4x-5\]Factor that and find out for what values of y is x=0.

OpenStudy (imstuck):

That factors very simply. Do you know what the factors are?

OpenStudy (imstuck):

And just so you know, this is a second degree polynomial which means it will have 2 zeros, not just one.